You are given a beaker of an unknown fluid, a beaker of water, and an irregular object of unknown density. The irregular object displaces 75% of its volume when floating in water, but 95% when it floats in the unknown liquid.

a) What is the (average) density of the irregular object?

b) What is the density of the unknown liquid?
Answer #1

relative density of the material of cube = Percentage of volumesubmerged)/100

RD = density of the object/density of water

density of water = 1 gm/cc

so

Desnity of irregular object = 75/100 = 0.75 * 1 = 0.75 gm/cc or 750 kg/m^3

--------------------------------------------


the relativedensity of the unknown fluid is RD = 75*100/95

RD = 0.789

Density D = 0.789*1 = 0.789 g/cc or 789 kg/m^3

------------------------------------